# Tensor Attention Training: Provably Efficient Learning of Higher-order Transformers

### v2

@@
Tensor Attention Training: Provably Efficient Learning of Higher-order Transformers
Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song magic.linuxkde@gmail.com. The Simons Institute for the Theory of Computing at the University of California, Berkeley.    Yufa Zhou yufazhou@seas.upenn.edu. University of Pennsylvania.
@@ 1 Introduction
The generative large language models (LLMs), such as Mistral [49], Llama [105], Llama2 [107], Llama3 [10], Gemma [106], GPT-3 [21], GPT-4 [2], Claude3 [12], Grok-1 [118] and many more, have been widely involved in people’s living and work in these two years, such as bio-informatics [108], coding [45], education [54], finance [72], law [100], and even writing NeurIPS conference reviews [59]. The success of LLMs is based on the transformer architecture introduced by [111], which also has been introduced into other modality [30], such as vision-language models, e.g., CLIP [87], Flamingo [7], LLaMA-Adapter [128, 36], LLava [68, 60], BLIP [69, 65], MiniGPT-4 [124], Qwen [17, 18], Gemini [103], MM1 [76].
The above open-sourced large models use two-view matrix attention, i.e., each attention score/entry is related to two tokens (one query token and one key token) to capture the data correlation. More specifically, let $Z$ be hidden representations and $Q=ZW_{Q},K=ZW_{K},V=ZW_{V}$ be the corresponding query, key, and value matrices after projections using weights $W_{Q},W_{K},W_{V}$ respectively. Then, the classical/matrix attention head can be written as $\mathsf{Att}(Z)=\mathsf{Softmax}(QK^{\top})V$.
@@
However, [96] theoretically and empirically shows that classical attention can capture pairwise correlation but not triple-wise correlation due to the representational limitations of matrix attention. In other words, one classical matrix attention head “cannot” capture the information relevant to multi-views simultaneously unless using multiple layers with careful architecture design. This poses a fundamental technical obstacle for multi-modality models to efficiently fuse multiple representations/views to capture the high-order correlation among them.
@@
2 Related Work
Fast attention computation.
In recent years, significant advances have been made in the development of efficient attention computation. One research direction involves employing low-rank approximations, polynomial kernel, or random features for the attention matrix [130, 1, 13, 52, 95, 4, 44, 113, 39, 40, 70, 102, 38], which scales the computational complexity sub-quadratically with sequence length. Another method explores patterns of sparse attention that lessen the computational load [93, 41, 63, 67, 64, 62]. Additionally, using linear attention as an alternative to softmax attention has emerged as a substantial area of study [104, 55, 97, 126, 94, 4, 101, 123, 66, 33]. These innovations have enhanced the capability of transformer-based models to handle longer sequences, thereby broadening their potential applications across various fields [27, 91, 86, 34, 80, 120, 9, 115, 16, 71, 23, 24, 47, 98].
Tensor computation for high-order representation.
Tensors excel over matrices in capturing higher-order relationships within data. Calculating low-rank factorizations or approximations of tensors is essential in a wide range of computer science applications, such as natural language processing [28, 74, 75, 22], computer vision [8, 85, 58, 25], computer graphics [112, 117, 110], security [3, 6, 51], and data mining [50, 89, 53, 79]. Moreover, tensors are crucial in numerous machine learning applications [84, 48, 42, 11, 129, 122, 99] and other diverse fields [90, 26, 125, 121, 88].
@@ 3 Preliminary
Basic notations. We use $[n]$ to denote $\{1,2,\dots,n\}$. We use $e_{i}$ to denote a column vector where only $i$-th location is $1$ and zeros everywhere else. We denote an all $1$ vector using ${\bf 1}_{n}\in\mathbb{R}^{n}$ We use $\langle a,b\rangle$ to denote the inner product of $a,b\in\mathbb{R}^{d}$ i.e. $\langle a,b\rangle:=\sum_{i=1}^{d}a_{i}b_{i}$. We use $\|x\|_{p}$ to denote the $\ell_{p}$ norm of a vector $x\in\mathbb{R}^{n}$, i.e. $\|x\|_{2}:=(\sum_{i=1}^{n}x_{i}^{2})^{1/2}$, and $\|x\|_{\infty}:=\max_{i\in[n]}|x_{i}|$. We use $\circ$ to denote the Hadamard product i.e. the $(i,j)$-entry of $A\circ B$ is $A_{i,j}B_{i,j}$.
We use $\operatorname{tr}[A]:=\sum_{i=1}^{n}A_{i,i}$ to denote the trace of a matrix $A\in\mathbb{R}^{n\times n}$. We use $\exp(A)$ to denote a matrix where $\exp(A)_{i,j}:=\exp(A_{i,j})$ for a matrix $A\in\mathbb{R}^{n\times d}$. We use $\|A\|_{\infty}$ to denote the $\ell_{\infty}$ norm of a matrix $A\in\mathbb{R}^{n\times d}$, i.e. $\|A\|_{\infty}:=\max_{i\in[n],j\in[d]}|A_{i,j}|$. We use $\|A\|_{F}$ to denote the Frobenius norm of a matrix $A\in\mathbb{R}^{n\times d}$, i.e. $\|A\|_{F}:=\sqrt{\sum_{i\in[n]}\sum_{j\in[d]}|A_{i,j}|^{2}}$. We use $\operatorname{poly}(n)$ to denote polynomial time complexity w.r.t $n$.
@@
Let $A\in\mathbb{R}^{n\times d}$. We use $a:=\operatorname{vec}(A)$ to denote a length $nd$ vector. We stack rows of $A$ into a column vector, i.e. $\operatorname{vec}(A):=[a_{1}^{\top},a_{2}^{\top},\dots,a_{n}^{\top}]^{\top}$ where $a_{i}^{\top}$ is the $i$-th row of $A$, or simply $\operatorname{vec}(A)_{j+(i-1)d}:=A_{i,j}$ for any $i\in[n],j\in[d]$, visualized in Fig. 2. Let $I_{d}\in\mathbb{R}^{d\times d}$ denote an identity matrix. Let $\mathsf{I}_{d}\in\mathbb{R}^{d\times d\times d}$ denote an identity tensor, i.e., the diagonal entries are $1$ and zeros everywhere else. Let $X\in\mathbb{R}^{d\times d^{2}}$. Let $x\in\mathbb{R}^{d^{3}}$ denote the vectorization of $X\in\mathbb{R}^{d\times d^{2}}$. Let $\mathsf{X}\in\mathbb{R}^{d\times d\times d}$ be the tesnorization of $X\in\mathbb{R}^{d\times d^{2}}$, where $\mathsf{X}_{a,b,c}=X_{a,(b-1)d+c}$ for any $a,b,c\in[d]$. And we have $X=\mathrm{mat}(\mathsf{X})$.
@@ 3.1 Definition of Tensor Operations
Given $K_{1}\in\mathbb{R}^{n_{1}\times d_{1}}$ and $K_{2}\in\mathbb{R}^{n_{2}\times d_{2}}$, for any $i_{1}\in[n_{1}],i_{2}\in[n_{2}],j_{1}\in[d_{1}],j_{2}\in[d_{2}]$, we define $K:=K_{1}\otimes K_{2}\in\mathbb{R}^{n_{1}n_{2}\times d_{1}d_{2}}$ as follows
@@
In this work, we will primarily use the following column-wise and row-wise versions of the Kronecker product, which are special kinds of Kronecker product.
@@
Given matrices $K_{1}\in\mathbb{R}^{n_{1}\times d},K_{2}\in\mathbb{R}^{n_{2}\times d}$, we define matrix $K:=K_{1}\oslash K_{2}\in\mathbb{R}^{n_{1}n_{2}\times d}$ as follows
$K_{i_{1}+(i_{2}-1)n_{1},j}:=(K_{1})_{i_{1},j}\cdot(K_{2})_{i_{2},j},~{}~{}~{}\forall i_{1}\in[n_{1}],i_{2}\in[n_{2}],j\in[d].$
Definition 3.3 ($\ominus$ row-wise Kronecker product).
Given matrices $K_{1}\in\mathbb{R}^{n\times d_{1}},K_{2}\in\mathbb{R}^{n\times d_{2}}$, we define matrix $K:=K_{1}\ominus K_{2}\in\mathbb{R}^{n\times d_{1}d_{2}}$ as follows
$K_{i,j_{1}+(j_{2}-1)d_{1}}:=(K_{1})_{i,j_{1}}\cdot(K_{2})_{i,j_{2}},~{}~{}~{}\forall i\in[n],j_{1}\in[d_{1}],j_{2}\in[d_{2}].$
@@ 3.2 Key Definitions of Tensor Attention
Then, based on Kronecker product, we define tensor attention in the following way.
@@
Given input matrices $Q,K_{1},K_{2},$ $V_{1},$ $V_{2}\in\mathbb{R}^{n\times d}$, compute the following matrix
@@
Our Definition 3.5 covers the self-attention setting, when the query/key/values $Q,K_{1},K_{2},V_{1},V_{2}$ follow Definition 3.4 where they share the same input. It is then a tensor self-attention, which can capture high-order information of the input $Z$. When the query/key/values have different inputs, it is then a tensor cross-attention that can capture high-order relationships among multiple inputs.
@@
For our focus, tensor attention training, we would like to find weights to fit the tensor attention to a desired output $E$. We first simplify the attention expression of Definition 3.5, whose inputs are from Definition 3.4 with weight matrices $W_{Q},W_{K_{1}},W_{K_{2}},W_{V_{1}},W_{V_{2}}\in\mathbb{R}^{d\times d}$. Let $X:=W_{Q}\cdot(W_{K_{1}}\oslash W_{K_{2}})^{\top}\in\mathbb{R}^{d\times d^{2}}$ and $Y:=W_{V_{1}}\oslash W_{V_{2}}\in\mathbb{R}^{d^{2}\times d}$. It can be verified that the tensor attention equals $D^{-1}\exp(ZX(Z\otimes Z)^{\top}/d)(Z\otimes Z)Y$. Then we model the tensor attention training as the following tensor attention optimization problem (where $A_{1},A_{2},A_{3},A_{4},A_{5}$ are introduced to replace $Z$ to capture more general settings such as cross-attention). See Figure 3 for an illustration.
@@
Suppose that $A_{1},A_{2},A_{3},A_{4},A_{5},E\in\mathbb{R}^{n\times d}$ and $Y_{1},Y_{2}\in\mathbb{R}^{d\times d}$ are given. We formulate the attention optimization problem as
$\min_{X\in\mathbb{R}^{d\times d^{2}}}\mathsf{Loss}(X):=0.5\|D(X)^{-1}\exp(A_{1}X(A_{2}\otimes A_{3})^{\top}/d)(A_{4}\otimes A_{5})Y-E\|_{F}^{2},$
@@
The naive gradient computation for the above tensor attention training takes $\Omega(n^{3})$ time. The gradient for $X$ is the bottleneck while that for $Y$ is not, since $A_{1}X(A_{2}\otimes A_{3})^{\top}\in\mathbb{R}^{n\times n^{2}}$ lies in the non-linear function $\mathsf{Softmax}$. Also, note that with gradients of $X$ and $Y$, it is easy to get the gradients of the weight matrices $W_{Q},W_{K_{1}},W_{K_{2}},W_{V_{1}},W_{V_{2}}$. Therefore, our main focus is the following Approximate Tensor Attention Loss Gradient Computation task.
@@
Let $\epsilon,B>0$. Let $A_{1},A_{2},A_{3},A_{4},A_{5},E\in\mathbb{R}^{n\times d}$ and let $X_{1},X_{2},X_{3},Y_{1},Y_{2}\in\mathbb{R}^{d\times d}$ (see Definition 3.8). Let $X=X_{1}\cdot(X_{2}\oslash X_{3})^{\top}\in\mathbb{R}^{d\times d^{2}}$. Assume that $\max\{\|A_{1}X_{1}\|_{\infty}$, $\|A_{2}X_{2}\|_{\infty}$, $\|A_{3}X_{3}\|_{\infty}$, $\|A_{4}Y_{1}\|_{\infty}$, $\|A_{5}Y_{2}\|_{\infty}\}\leq B$. Let us assume that any numbers in the previous matrices are in the $\log(n)$ bits model. We define $\mathsf{Loss}(X)$ the same as Definition 3.8. Let the gradient of loss function $\mathsf{Loss}(X)$ be $\frac{\mathrm{d}\mathsf{Loss}(X)}{\mathrm{d}X}\in\mathbb{R}^{d\times d^{2}}$. Then, our target is to output a matrix $\widetilde{g}\in\mathbb{R}^{d\times d^{2}}$ satisfying $\|\widetilde{g}-\frac{\mathrm{d}\mathsf{Loss}(X)}{\mathrm{d}X}\|_{\infty}\leq\epsilon.$

### v3

# Training Tensor Attention Efficiently: From Cubic to Almost Linear Time

@@
Training Tensor Attention Efficiently: From Cubic to Almost Linear Time
Yang Cao and Yingyu Liang and Zhenmei Shi and Zhao Song Wyoming Seminary.The University of Hong Kong.University of Wisconsin-Madison.magic.linuxkde@gmail.com. Simons Institute for the Theory of Computing, University of California, Berkeley.
@@ 1 Introduction
The generative large language models (LLMs), such as Mistral 3.1 [63], Llama 4 [61], Gemma 3 [82], GPT-4o [66], Claude 3.7 [9], Grok 3 [91], Qwen 3 [69], DeepSeek R1 [31], and many more have been widely involved in people’s living and work in these two years, such as bio-informatics [83], coding [34], education [39], finance [49], law [80], and even writing NeurIPS conference reviews [50]. The success of LLMs is based on the transformer architecture introduced by [87], which also has been introduced into other modalities [24], such as vision-language models, e.g., CLIP [70], Flamingo [4], LLaMA-Adapter [97, 27], LLava [53, 52], BLIP [48, 47], MiniGPT-4 [102], Qwen [10, 11], Gemini [81], MM1 [59].
The above open-sourced large models use two-view matrix attention, i.e., each attention score/entry is related to two tokens (one query token and one key token) to capture the data correlation. More specifically, let $Z$ be hidden representations and $Q=ZW_{Q},K=ZW_{K},V=ZW_{V}$ be the corresponding query, key, and value matrices after projections using weights $W_{Q},W_{K},W_{V}$, respectively. Then, the classical/matrix attention head can be written as $\mathsf{Att}(Z)=\mathsf{Softmax}(QK^{\top})V$.
@@
However, [74] theoretically and empirically shows that classical attention can capture pairwise correlation but not triple-wise correlation due to the representational limitations of matrix attention. [74] introduces a triple detection task, demonstrating that classical attention layers have complexity scaling linearly with the input size, while high-order attention can efficiently perform this computation within a single head.
In other words, one classical matrix attention head “cannot” capture the information relevant to multi-views simultaneously unless using multiple layers with careful architecture design. This poses a fundamental technical obstacle for multi-modality models to efficiently fuse multiple representations/views to capture the high-order correlation among them, e.g., the high-order correlations among multi-modalities such as audio, text, and images.
@@ 2 Preliminary
Basic notations.
We use $[n]$ to denote $\{1,2,\dots,n\}$. We use $e_{i}$ to denote a column vector where only $i$-th location is $1$ and zeros everywhere else. We denote an all $1$ vector using ${\bf 1}_{n}\in\mathbb{R}^{n}$ . We use $\langle a,b\rangle$ to denote the inner product of $a,b\in\mathbb{R}^{d}$ i.e. $\langle a,b\rangle:=\sum_{i=1}^{d}a_{i}b_{i}$. We use $\|x\|_{p}$ to denote the $\ell_{p}$ norm of a vector $x\in\mathbb{R}^{n}$, i.e. $\|x\|_{2}:=(\sum_{i=1}^{n}x_{i}^{2})^{1/2}$, and $\|x\|_{\infty}:=\max_{i\in[n]}|x_{i}|$. We use $\circ$ to denote the Hadamard product, i.e., the $(i,j)$-entry of $A\circ B$ is $A_{i,j}B_{i,j}$. We use $\operatorname{tr}[A]:=\sum_{i=1}^{n}A_{i,i}$ to denote the trace of a matrix $A\in\mathbb{R}^{n\times n}$. We use $\exp(A)$ to denote a matrix where $\exp(A)_{i,j}:=\exp(A_{i,j})$ for a matrix $A\in\mathbb{R}^{n\times d}$. We use $\|A\|_{\infty}$ to denote the $\ell_{\infty}$ norm of a matrix $A\in\mathbb{R}^{n\times d}$, i.e. $\|A\|_{\infty}:=\max_{i\in[n],j\in[d]}|A_{i,j}|$. We use $\|A\|_{F}$ to denote the Frobenius norm of a matrix $A\in\mathbb{R}^{n\times d}$, i.e. $\|A\|_{F}:=\sqrt{\sum_{i\in[n]}\sum_{j\in[d]}|A_{i,j}|^{2}}$. We use $\operatorname{poly}(n)$ to denote polynomial time complexity w.r.t. $n$.
@@
Let $A\in\mathbb{R}^{n\times d}$. We use $a:=\operatorname{vec}(A)$ to denote the length $nd$ vector obtained by stacking rows of $A$ into a column vector, i.e. $\operatorname{vec}(A):=[a_{1}^{\top},a_{2}^{\top},\dots,a_{n}^{\top}]^{\top}$ where $a_{i}^{\top}$ is the $i$-th row of $A$, or simply $\operatorname{vec}(A)_{j+(i-1)d}:=A_{i,j}$ for any $i\in[n],j\in[d]$, visualized in Fig. 2. Let $I_{d}\in\mathbb{R}^{d\times d}$ denote the identity matrix. Let $\mathsf{I}_{d}\in\mathbb{R}^{d\times d\times d}$ denote the identity tensor, i.e., the diagonal entries are $1$ and zeros everywhere else. Let $X\in\mathbb{R}^{d\times d^{2}}$. Let $x\in\mathbb{R}^{d^{3}}$ denote the vectorization of $X\in\mathbb{R}^{d\times d^{2}}$. Let $\mathsf{X}\in\mathbb{R}^{d\times d\times d}$ be the tensorization of $X\in\mathbb{R}^{d\times d^{2}}$, where $\mathsf{X}_{a,b,c}=X_{a,(b-1)d+c}$ for any $a,b,c\in[d]$. We define the corresponding function $\mathrm{mat}:\mathbb{R}^{d\times d\times d}\to\mathbb{R}^{d\times d^{2}}$ as $X=\mathrm{mat}(\mathsf{X})$ where $X_{a,(b-1)d+c}=\mathsf{X}_{a,b,c}$ for any $a,b,c\in[d]$.
@@ 2.1 Definition of Tensor Operations
Given $K_{1}\in\mathbb{R}^{n_{1}\times d_{1}}$ and $K_{2}\in\mathbb{R}^{n_{2}\times d_{2}}$, for any $i_{1}\in[n_{1}],i_{2}\in[n_{2}],j_{1}\in[d_{1}],j_{2}\in[d_{2}]$, we define the matrix $K:=K_{1}\otimes K_{2}\in\mathbb{R}^{n_{1}n_{2}\times d_{1}d_{2}}$ as
@@
In this work, we will primarily use the following column-wise and row-wise versions of the Kronecker product, which are special kinds of Kronecker products.
@@
Given matrices $K_{1}\in\mathbb{R}^{n_{1}\times d},K_{2}\in\mathbb{R}^{n_{2}\times d}$, we define the matrix $K:=K_{1}\oslash K_{2}\in\mathbb{R}^{n_{1}n_{2}\times d}$ as follows
$K_{i_{1}+(i_{2}-1)n_{1},j}:=(K_{1})_{i_{1},j}\cdot(K_{2})_{i_{2},j},$
$\forall i_{1}\in[n_{1}],i_{2}\in[n_{2}],j\in[d].$
Definition 2.3 ($\ominus$ row-wise Kronecker product, also referred to as the face-splitting product).
Given matrices $K_{1}\in\mathbb{R}^{n\times d_{1}},K_{2}\in\mathbb{R}^{n\times d_{2}}$, we define the matrix $K:=K_{1}\ominus K_{2}\in\mathbb{R}^{n\times d_{1}d_{2}}$ as follows
$K_{i,j_{1}+(j_{2}-1)d_{1}}:=(K_{1})_{i,j_{1}}\cdot(K_{2})_{i,j_{2}},$
$\forall i\in[n],j_{1}\in[d_{1}],j_{2}\in[d_{2}]$.
@@ 2.2 Key Definitions of Tensor Attention
Then, based on the Kronecker product, we define tensor attention in the following way.
@@
Given input matrices $Q,K_{1},K_{2},$ $V_{1},$ $V_{2}\in\mathbb{R}^{n\times d}$, compute the matrix
@@
Our Definition 2.5 covers the self-attention setting when the query/key/values $Q,K_{1},K_{2},V_{1},V_{2}$ follow Definition 2.4 where they share the same input. It is then a tensor self-attention, which can capture high-order information of the input $Z$. When the query/key/values have different inputs, it is then a tensor cross-attention that can capture high-order relationships among multiple inputs.
@@
For our focus, tensor attention training, we would like to find weights to fit the tensor attention to a desired output $E$. We first simplify the attention expression of Definition 2.5, whose inputs are from Definition 2.4 with weight matrices $W_{Q},W_{K_{1}},W_{K_{2}},W_{V_{1}},W_{V_{2}}\in\mathbb{R}^{d\times d}$. Let $X:=W_{Q}\cdot(W_{K_{1}}\oslash W_{K_{2}})^{\top}\in\mathbb{R}^{d\times d^{2}}$ and $Y:=W_{V_{1}}\oslash W_{V_{2}}\in\mathbb{R}^{d^{2}\times d}$. It can be verified that the tensor attention equals
$D^{-1}\exp(ZX(Z\otimes Z)^{\top}/d)(Z\otimes Z)Y,$
where $Z\in\mathbb{R}^{n\times d}$ is defined as the input sequence in Definition 2.4.
The naive gradient computation for the tensor attention training takes $\Omega(n^{3})$ time. The gradient for $X$ is the bottleneck while that for $Y$ is not, since $A_{1}X(A_{2}\otimes A_{3})^{\top}\in\mathbb{R}^{n\times n^{2}}$ lies in the non-linear function $\mathsf{Softmax}$. Also, note that with gradients of $X$ and $Y$, it is easy to get the gradients of the weight matrices $W_{Q},W_{K_{1}},W_{K_{2}},W_{V_{1}},W_{V_{2}}$. Therefore, we model the tensor attention training as the following tensor attention optimization problem (where $A_{1},A_{2},A_{3},A_{4},A_{5}$ are introduced to replace $Z$ to capture more general settings such as cross-attention). See Figure 3 for an illustration.
@@
Suppose that $A_{1},A_{2},A_{3},A_{4},A_{5},E\in\mathbb{R}^{n\times d}$ and $Y_{1},Y_{2}\in\mathbb{R}^{d\times d}$ are given. We formulate the attention optimization problem
$\min_{X\in\mathbb{R}^{d\times d^{2}}}\mathsf{Loss}(X)$
as
$0.5\|D(X)^{-1}\exp(A_{1}X(A_{2}\otimes A_{3})^{\top}/d)(A_{4}\otimes A_{5})Y-E\|_{F}^{2},$
@@
Our main focus is the following Approximate Tensor Attention Loss Gradient Computation task.
@@
Let $\epsilon,B>0$. Let $A_{1},A_{2},A_{3},A_{4},A_{5},E\in\mathbb{R}^{n\times d}$ and let $X_{1},X_{2},X_{3},Y_{1},Y_{2}\in\mathbb{R}^{d\times d}$ (see Definition 2.8). Let $X=X_{1}\cdot(X_{2}\oslash X_{3})^{\top}\in\mathbb{R}^{d\times d^{2}}$. Assume that $\max\{\|A_{1}X_{1}\|_{\infty}$, $\|A_{2}X_{2}\|_{\infty}$, $\|A_{3}X_{3}\|_{\infty}$, $\|A_{4}Y_{1}\|_{\infty}$, $\|A_{5}Y_{2}\|_{\infty}\}\leq B$. Let us assume that any numbers in the previous matrices are in the $\log(n)$ bits model [^1] Each entry in the matrix is represented by at most $\log(n)$ bits. This assumption is well-accepted and widely used in the computational complexity community, e.g., [26, 51, 60]. . We define $\mathsf{Loss}(X)$ the same as Definition 2.8. Let the gradient of loss function $\mathsf{Loss}(X)$ be $\frac{\mathrm{d}\mathsf{Loss}(X)}{\mathrm{d}X}\in\mathbb{R}^{d\times d^{2}}$. Then, our target is to output a matrix $\widetilde{g}\in\mathbb{R}^{d\times d^{2}}$ satisfying
$\|\widetilde{g}-\frac{\mathrm{d}\mathsf{Loss}(X)}{\mathrm{d}X}\|_{\infty}\leq\epsilon.$
Figure 4: The computational graph for tensor attention backward. The blue boxes are input matrices, the gray boxes are intermediate matrices, and the orange box is the final gradient matrix. Here, $A_{1},A_{2},A_{3},A_{4},A_{5}$ denote the previous inputs, $E$ denotes the target matrix, and $X,Y$ denote the attention weights. More detailed definitions of each variable can be found in Section D, E and F.